X^2p^2+xyp+2y^2=0
1
Expert's answer
2020-10-21T17:12:31-0400

Given

"x^2p^2+xyp+2y^2=0"

Let us denote "xp=q" ,thus we get "q^2+yq+2y^2=0" , now solving q ,we get,


"q=\frac{-y\pm\sqrt{y^2-4\cdot 1\cdot 2y^2}}{2}\\
\implies q=\frac{-y\pm y\sqrt{7}i}{2}\\
\implies xp=x\frac{dy}{dx}=y(-1\pm\sqrt{7}i)\\
\implies \frac{dy}{y}=(-1\pm\sqrt{7}i)\frac{dx}{x}\\
\implies \int\frac{dy}{y}=(-1\pm\sqrt{7}i)\int\frac{dx}{x}\\
\implies \ln y=-1\pm\sqrt{7}i\ln x+\ln c\\
\implies y=cx^{-1\pm\sqrt{7}i}"

Where c is constant and "i=\sqrt{-1}"